OpenStudy (anonymous):

Simplify the expression 83 + 263 + 17 - 63 A. 426 B. 300 C. 306 D. 200

OpenStudy (anonymous):

B. 300

OpenStudy (anonymous):

This is one you might easily check on a calculator. If you're trying to do it quickly, though, look for terms that, when added together, make the whole calculation easier. No need to just add it all up at once. Like 263-63 = 200. 83+17 = 100. 100+200 = 300.